OpenStudy (anonymous):

What is the simplified form of each expression?? (-2x^8)*3y^9*2x^4

OpenStudy (anonymous):

Will give medal

OpenStudy (anonymous):

@ZeHanz can you help im super lost

OpenStudy (zehanz):

I think your expression is: \((-2x^{8})\cdot 3y^{9}\cdot 2x^{4}\). You can rearrange the factors: \(-2\cdot 3\cdot 2\cdot x^{8}\cdot x^{4}\cdot y^{9}\) Now use the power rule \(x^{n}\cdot x^{m}=x^{n+m}\). Can you do that?

OpenStudy (anonymous):

I think so but i can see the numbers to the power it looks like x^72*x^192=x^72+m are those the right numbers?

OpenStudy (anonymous):

would the answer be 1.5x10^12?

OpenStudy (zehanz):

That doesn't make sense. 1.5x10^12 = 1,500,000,000,000, so that is just a very large number. In the expression you gave, there are x's and y's. After simplifying the expression, there will still be x's and y's. My answer would be: \(-12x^{12}y^{9}\)

OpenStudy (anonymous):

mmmm thats true but all of my answer choices dont come close to yours?

OpenStudy (zehanz):

Can you show a picture of the expression, to be sure what it is?